# Visual Merit or Linguistic Crutch? A Close Look at DeepSeek-OCR

### v1

@@ Abstract
Yunhao Liang1,2thanks:  Equal Contribution.    Ruixuan Ying311footnotemark: 1    Bo Li411footnotemark: 1   Hong Li4   Kai Yan4   Qingwen Li4 Min Yang7,8   Okamoto Satoshi3,4,5   Zhe Cui1,2   Shiwen Ni7,8thanks:  Corresponding author. 🖂 sw.ni@siat.ac.cn 1Chengdu Institute of Computer Applications, Chinese Academy of Sciences, Chengdu, China 2University of Chinese Academy of Sciences, Beijing, China 3Institute of Multidisciplinary Research for Advanced Materials (IMRAM), Tohoku University, Sendai, Japan 4China Tower Corporation Limited, Beijing, China 5Center for Science and Innovation in Spintronics (CSIS), Tohoku University, Sendai, Japan 6National Institute for Materials Science (NIMS), Tsukuba, Japan 7Shenzhen Key Laboratory for High Performance Data Mining, Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences, Shenzhen, China 8Artificial Intelligence Research Institute, Shenzhen University of Advanced Technology, Shenzhen, China

### v2

@@ Abstract
Yunhao Liang1,2thanks:  Equal Contribution.    Ruixuan Ying311footnotemark: 1    Bo Li4   Hong Li4   Kai Yan4   Feiteng Fang7 Qingwen Li4   Min Yang7,8   Okamoto Satoshi3,5,6   Zhe Cui1,2   Shiwen Ni7,8thanks:  Corresponding author. 🖂 sw.ni@siat.ac.cn 1Chengdu Institute of Computer Applications, CAS   2University of Chinese Academy of Sciences 3IMRAM, Tohoku University  4China Tower Corporation Limited  5CSIS, Tohoku University 6National Institute for Materials Science  7Shenzhen Institutes of Advanced Technology, CAS 8Artificial Intelligence Research Institute, Shenzhen University of Advanced Technology